The phase difference between two points separated by 1 m in a wave of frequency 120hz is 90 degree .The wave velocity will be …....?

Grade:12th pass

2 Answers

Arun
6 years ago
Dear Gaurav
 
Path difference \Delta = \lambda \phi/2\pi
 
1 = (\lambda/2\pi) * (\pi/2)
\lambda = 4 m
 
Hence
V = n* \lambda = 120*4 = 480 m/sec.
 
Regards
Arun (askIITians forum expert)
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Path difference Δ = λɸ/2π
1 = (λ/2π)*(π/2)
λ = 4 m
Hence
V = n*λ = 120*4 = 480 m/sec.

Thanks and Regards